# Towards Competitive Search Relevance For Inference-Free Learned Sparse Retrievers

### v1

@@ Abstract.
Learned sparse retrieval, which can efficiently perform retrieval through mature inverted-index engines, has garnered growing attention in recent years. Particularly, the inference-free sparse retrievers are attractive as they eliminate online model inference in the retrieval phase thereby avoids huge computational cost, offering reasonable throughput and latency. However, even the state-of-the-art (SOTA) inference-free sparse models lag far behind in terms of search relevance when compared to both sparse and dense siamese models. Towards competitive search relevance for inference-free sparse retrievers, we argue that they deserve dedicated training methods other than using same ones with siamese encoders. In this paper, we propose two different approaches for performance improvement. First, we introduce the IDF-aware FLOPS loss, which introduces Inverted Document Frequency (IDF) to the sparsification of representations. We find that it mitigates the negative impact of the FLOPS regularization on search relevance, allowing the model to achieve a better balance between accuracy and efficiency. Moreover, we propose a heterogeneous ensemble knowledge distillation framework that combines siamese dense and sparse retrievers to generate supervisory signals during the pre-training phase. The ensemble framework of dense and sparse retriever capitalizes on their strengths respectively, providing a strong upper bound for knowledge distillation. To concur the diverse feedback from heterogeneous supervisors, we normalize and then aggregate the outputs of the teacher models to eliminate score scale differences. On the BEIR benchmark, our model outperforms existing SOTA inference-free sparse model by 3.3 NDCG@10 score. It exhibits search relevance comparable to siamese sparse retrievers and client-side latency only 1.1x that of BM25.
@@ 1. Introduction
In this paper, we focus on improving the search relevance of the inference-free sparse retriever through more effective training methodologies. The first challenge lies in the uniform penalty applied to all tokens by the FLOPS regularization (Paria et al., 2020).
The purpose of the FLOPS loss is to eliminate unnecessary tokens to sparsify the document representation. However, the uniform penalty from FLOPS loss has a side effect: it prevents the model from assigning tokens with large weights. Consequently, for semantically rich tokens that are important to be retained in the representation, their importance will be underestimated. To mitigate this effect, we propose a simple yet effective enhancement called IDF-aware FLOPS. This loss adjusts the penalty imposed by the FLOPS loss on different tokens, making the penalty inversely proportional to the token’s IDF value. It reduces the regularization penalty on unique tokens and increases the penalty on frequent tokens.
Through experiments, we demonstrate that IDF-aware FLOPS effectively improves the search relevance of the inference-free sparse retriever, and we discover that it effectively reduces the average FLOPS number for the retrieval process.
@@
We conduct experiments on 13 public datasets from the BEIR benchmark, and our model outperforms the existing SOTA inference-free sparse model by 3.3 average NDCG@10 scores. Its performance even surpasses many strong siamese retrievers. Our contributions can be summarized as follows: (1) We propose IDF-aware FLOPS, which effectively improves the search relevance and efficiency of inference-free sparse models. (2) We explore how to effectively pre-train inference-free sparse models and propose the ensemble teacher model of heterogeneous siamese models, which has reasonable inference costs and strong performance. (3) The zero-shot performance of our model outperforms the SOTA inference-free retriever by 3.3 NDCG@10 score. It also surpasses strong siamese retrievers including SPLADE-v3-DistilBERT and ColBERTv2. While its client-side latency is only 1.1x that of BM25.
@@ 3. Preliminary
3.1. Model
SPLADE-doc-distill is an expansion model that predicts the estimated importance of tokens in the vocabulary space. The model reuses the masked language model head to generate a similarity distribution over the token vocabulary. Utilizing BERT as its backbone, the model operates with a sparse vector of 30,522 dimensions. It utilizes max pooling to aggregate the predictions across different positions and uses ReLU activation to sparsify the representation and ensure the weights are non-negative. SPLADE-doc-distill has an asymmetric architecture, which means it only performs model inference on documents. For an input document, let $w_{i,j}$ represent the weight of token $j$ at position $i$. The formula to get the sparse representation is as follows:
(1) $w_{j}=\max_{i}\log(1+\text{ReLU}(w_{i,j}))$
The sparsity arises from the FLOPs regularizer during training and the ReLU activation. For query representation, it can be illustrated as binary Bag-of-Words. That is, the weight $w_{j}$ of token $j$ in query representation is set to 1 only if the token $j$ is present in the query; otherwise, $w_{j}$ is set to 0. Consequently, the ranking score for a given query and document can be computed as the inner product of their sparse representations, which is formulated as follows:
(2) $s(q,d)=\sum_{j\in q}w_{j}^{d}$
3.2. Training
The optimization objective of SPLADE combines the ranking loss $\mathcal{L}_{\text{rank}}$ and sparse regularization loss $\mathcal{L}^{d}_{\texttt{FLOPS}}$, ensuring an end-to-end training approach that simultaneously optimizes the search relevance and representation sparsification. For SPLADE-doc-distill, the training loss $\mathcal{L}$ is formulated as follows:
(3) $\mathcal{L}=\mathcal{L}_{\texttt{rank}}+\lambda_{d}\mathcal{L}^{d}_{\texttt{FLOPS}}$
(4) $\mathcal{L}^{d}_{\texttt{FLOPS}}=\sum_{j\in V}{\bar{a}}_{j}^{2}=\sum_{j\in V}\left(\frac{1}{N}\sum_{i=1}^{N}w_{j}^{(d_{i})}\right)^{2}$
where $\lambda_{d}$ is the weight for FLOPS regularizer and $\overline{a}_{j}$ is the average weight of token $j$ in the batch. The FLOPS regularizer penalizes the dimensions with high average weights in the representation to achieve the sparsity. In the context of text retrieval, tokens occurs more frequently and tokens with larger weights get more penalty. For given triplets $(q,d^{+},d^{-})$ consisting of a query, a positive document, and a hard negative document, SPLADE-doc-distill utilizes the KL divergence function to fit the match score differences produced by the model on the given triplets to the predicted results of the cross-encoder teacher model, thereby performing knowledge distillation.
@@ 4. Method
4.1. IDF-aware FLOPS
For the vanilla SPLADE-doc-distill model, according to Equation 4, we note that the FLOPS loss imposes a uniform penalty on different tokens. This penalty increases in a quadratic way with the average weights of token $j$, while the growth rate of the match score on token $j$ is linear. Although the initial intention of the FLOPS regularizer was to sparsify the representation, one unintended consequence is that it prevents the model from generating excessively large weight for a single token. For semantically important tokens that are always retained once they occurs, FLOPS loss put an extra restriction on their weights.
Therefore, we propose the IDF-aware FLOPS, which is a simple yet highly effective approach. We multiply the model output with token IDF values as the token weight. When computing the FLOPS loss, we exclude the IDF values and only apply constraints on the model output. The IDF values are involved in the calculation of the match score during both training and inference phases. Formally, we modify the token weight and FLOPS loss formulas for Equation 1 and Equation 4 to the following:
(5) $w_{j}=\text{IDF}_{j}\cdot\max_{i}\log(1+\text{ReLU}(w_{i,j}))$
(6) $\mathcal{L}^{d}_{\texttt{FLOPS}}=\sum_{j\in V}{\bar{a}}_{j}^{2}=\sum_{j\in V}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{w_{j}^{(d_{i})}}{\text{IDF}_{j}}\right)^{2}$
where the other formulas and procedures remain unchanged.
This approach ensures that the penalty imposed by the FLOPS regularization on the model outputs is inversely proportional to the token’s IDF value. This compensates for the potential underestimation of important tokens. Furthermore, IDF-aware FLOPS drastically increases the penalty on low-IDF tokens, efficiently reducing the complexity of the retrieval process.
To present a straightforward view on the typical low-IDF and high-IDF tokens, we conduct a case study on the MS MARCO dataset [^2] https://github.com/microsoft/MSMARCO-Passage-Ranking. We utilize the uncased BERT tokenizer, and the outcome is presented in Table 1. According to the results, tokens with the highest IDF values represent unique semantics. We should always keep them by giving less penalty. In contrast, tokens with the lowest IDF values are more akin to stopwords. Imposing a uniform regularization penalty on these tokens would be an unreasonable approach.
Table 1. Case study of IDF values and tokens on the MS MARCO dataset. We only consider tokens appearing in more than 1000 documents.
Lowest IDF Token IDF Value Highest IDF Token IDF Value
. 0.016 janeiro 9.086
the 0.135 yemen 9.086
, 0.176 interiors 9.086
of 0.298 underwood 9.086
and 0.318 trevor 9.085
As a widely subscribed factor in relevance calculation (such as TF-IDF or BM25), the IDF values are tracked real-time in most of the lexical search engines. And this provides the convenience of obtaining the IDF values. Furthermore, We find that precomputed IDF values derived from the training set also exhibit strong generalization capabilities. Details can be found in Section 5.4.1.
The standard SPLADE-doc model assigns a weight of 1 to all query tokens. By decoupling the IDF values and the model outputs, we can utilize the IDF values as query token weights when constructing the query body in the search engine. This enables the application of heuristic optimization rules, further enhancing search latency and throughput. Details can be found in Section 5.3.2.
@@ 5.1.2. Model Training
For the teacher models of the pre-training phase, we employ SOTA dense and sparse retrievers , namely gte-large-en-v1.5666https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5 and opensearch-neural-sparse-encoding-v1777https://huggingface.co/opensearch-project/opensearch-neural-sparse-encoding-v1. The total number of pre-training steps is 150,000. For each step, we sample 48 queries and 8 hard samples(1 positive document and 7 hard negative documents) for every query. We set the learning rate to 5e-5 and the FLOPS $\lambda_{d}$ to 1e-7. The max input length is set to 128. The constant $S$ is set to 10 for scaling back the assembled scores. In the fine-tuning phase, we utilize an ensemble teacher model comprising the two siamese retrievers used in the pre-training phase, as well as two cross-encoder re-rankers [^8] We’re using https://huggingface.co/castorini/monot5-3b-msmarco-10k and https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-12-v2 . The total number of training steps is 50,000. For each step, we sample 40 queries and 10 hard negatives for each query. We set the learning rate to 2e-5 and the FLOPS $\lambda_{d}$ to 0.02. The max input length is set to 256. And the IDF values are calculated using the documents of MS MARCO dataset. The constant $S$ is set to 30 for scaling back the assembled scores.
@@ 5.4.1. Impact on search relevance
To assess the impact of IDF-aware FLOPS on the zero-shot search relevance, we conduct an ablation study employing different IDF settings on BEIR benchmark. In the default setting, we use fixed IDF values derived from MS MARCO dataset. We also conduct experiments using IDF values from the corresponding BEIR datasets. With different settings, we examined the impact of (1) employing IDF-aware FLOPS in training phase (2) retrieval with IDF values derived from different sources. We conduct these experiments on our model and SPLADE-doc-distill. The experiment results are shown in Table 6. From the experiment results, we obtain several conclusions: (1) The IDF-aware FLOPS boost the model search relevance at large margin. Training and inference with IDF, both our model and SPLADE-doc-distill achieve much better search relevance. (2) For models without pre-training phase, using IDF derived from the test set has better search relevance compared with fixed IDF derived from training data. However, if the model has undergone extensive pre-training on large-scale data using the fixed IDF, the conclusion is the opposite.
@@ 5.4.2. Impact on retrieval efficiency and index size
Since the IDF-aware FLOPS applies larger penalty on low IDF tokens, experiments are conducted to measure its impact on retrieval efficiency. We conducted experiments to quantify the relationship between the expansion rate(average token number per expanded documents), retrieval efficiency(FLOPS number), and search relevance for our model. The comparison is made between our model with IDF-aware FLOPS and our model trained without IDF-aware FLOPS. The experimental results are depicted in Figure 3. The experimental findings demonstrate that the IDF-aware FLOPS substantially augments the retrieval efficiency. For models with similar expansion rate, the FLOPS number of those trained with IDF-aware FLOPS is much smaller.
Figure 3. The average FLOPS number, average token number per document and search relevance on BEIR datasets for models with and without IDF-aware FLOPS. The models are trained with different $\lambda_{d}$.
@@ 6. Conclusion
In this paper, we proposed two novel approaches to significantly improve the search relevance of inference-free learned sparse retrievers while maintaining high efficiency. We introduced IDF-aware FLOPS to mitigate underestimation of important tokens and increase sparsity, boosting both relevance and efficiency. We also developed a heterogeneous ensemble knowledge distillation framework leveraging strong dense and sparse retrievers for pre-training, enhancing generalization. Extensive experiments validate our methods’ effectiveness. Our model achieves SOTA performance among inference-free sparse retrievers on BEIR, while maintaining end-to-end latency only 1.1x that of BM25.

### v2

@@ Abstract.
Learned sparse retrieval, which can efficiently perform retrieval through mature inverted-index engines, has garnered growing attention in recent years. Particularly, the inference-free sparse retrievers are attractive as they eliminate online model inference in the retrieval phase thereby avoids huge computational cost, offering reasonable throughput and latency. However, even the state-of-the-art (SOTA) inference-free sparse models lag far behind in terms of search relevance when compared to both sparse and dense siamese models. Towards competitive search relevance for inference-free sparse retrievers, we argue that they deserve dedicated training methods other than using same ones with siamese encoders. In this paper, we propose two different approaches for performance improvement. First, we propose an IDF-aware penalty for the matching function that suppresses the contribution of low-IDF tokens and increases the model’s focus on informative terms. Moreover, we propose a heterogeneous ensemble knowledge distillation framework that combines siamese dense and sparse retrievers to generate supervisory signals during the pre-training phase. The ensemble framework of dense and sparse retriever capitalizes on their strengths respectively, providing a strong upper bound for knowledge distillation. To concur the diverse feedback from heterogeneous supervisors, we normalize and then aggregate the outputs of the teacher models to eliminate score scale differences. On the BEIR benchmark, our model outperforms existing SOTA inference-free sparse model by 3.3 NDCG@10 score. It exhibits search relevance comparable to siamese sparse retrievers and client-side latency only 1.1x that of BM25.
@@ 1. Introduction
In this paper, we focus on improving the search relevance of the inference-free sparse retriever through more effective training methodologies. The first challenge lies in the uniform penalty applied to all tokens by the FLOPS regularization (Paria et al., 2020). We argue that the uniform penalty unintentionally suppresses rare but important terms. To address this, we propose a penalty strategy that adjusts token importance based on their inverse document frequency (IDF). Specifically, we integrate IDF weights into the scoring mechanism, which encourages the model to assign higher relevance scores to informative low-frequency tokens. This, in turn, guides the optimization to preserve such tokens during training, improving relevance and maintaining overall sparsity. Through experiments, we demonstrate that IDF-aware penalty effectively improves the search relevance of the inference-free sparse retriever, and we discover that it effectively reduces the average FLOPS number for the retrieval process.
@@
We conduct experiments on 13 public datasets from the BEIR benchmark, and our model outperforms the existing SOTA inference-free sparse model by 3.3 average NDCG@10 scores. Its performance even surpasses many strong siamese retrievers. Our contributions can be summarized as follows: (1) We propose IDF-aware penalty, which effectively improves the search relevance and efficiency of inference-free sparse models. (2) We explore how to effectively pre-train inference-free sparse models and propose the ensemble teacher model of heterogeneous siamese models, which has reasonable inference costs and strong performance. (3) The zero-shot performance of our model outperforms the SOTA inference-free retriever by 3.3 NDCG@10 score. It also surpasses strong siamese retrievers including SPLADE-v3-DistilBERT and ColBERTv2. While its client-side latency is only 1.1x that of BM25.
@@ 3. Preliminary
3.1. Ranking supervision.
To capture semantic relevance, SPLADE-doc-distill employs knowledge distillation in which the student model learns from the teacher models. Let $\mathbf{s}_{\text{tea}}$ and $\mathbf{s}_{\text{stu}}$ denote the teacher and student scores across documents. The ranking loss is defined as the KL divergence between their softmax distributions:
(1) $\mathcal{L}_{\text{rank}}=\text{KL}\big{(}\,\text{softmax}(\mathbf{s}_{\text{tea}})\|\,\text{softmax}(\mathbf{s}_{\text{stu}})\,\big{)},$
where $\mathbf{s}_{\text{stu}}=[s(q,d_{1}),\dots,s(q,d_{n})]$ denotes the scores computed by the student model between the query $q$ and each candidate document $d_{i}$, and
(2) $s(q,d_{i})=\sum_{t\in\mathcal{V}}q_{t}\cdot d_{i,t},$
with $q_{t}\in\{0,1\}$ indicating the presence of token $t$ in query $q$, and $d_{i,t}$ representing the activation value of token $t$ in document $d_{i}$.
3.2. Sparsity regularization.
To promote sparsity in the document vectors, it applies a FLOPS regularizer (Paria et al., 2020) that penalizes high average token activations. For a batch of $N$ documents, let $w_{j}^{(d_{i})}$ be the activation of token $j$ in document $d_{i}$. The FLOPS loss is defined as the sum of squared average activations:
(3) $\mathcal{L}_{\text{FLOPS}}=\sum_{j\in\mathcal{V}}\left(\frac{1}{N}\sum_{i=1}^{N}w_{j}^{(d_{i})}\right)^{2}.$
The final training loss jointly optimizes relevance alignment and representation sparsity:
(4) $\mathcal{L}=\mathcal{L}_{\text{rank}}+\lambda_{d}\cdot\mathcal{L}_{\text{FLOPS}},$
where $\lambda_{d}$ is a hyperparameter controlling the trade-off between matching accuracy and sparsity.
@@ 4. Method
4.1. IDF-aware Penalty
Observation.
To better understand how existing sparsity regularization behaves in retrieval models, we analyze token-level activation and FLOPs penalty scores using a standard SPLADE-v3-doc model (Lassance et al., 2024) on scidocs dataset (Thakur et al., 2021).
Figure 1. Token-level activation vs. FLOPs penalty in a random batch encoded by SPLADE-v3-doc. The sparsity regularization under-penalizes trivial tokens while suppresses informative ones.
As illustrated in Figure 1, the distribution shows that many trivial tokens—such as ”that” and ”this”—receive low FLOPs penalties despite being activated, while more meaningful tokens like ”algorithms” or ”graph” are penalized disproportionately. This indicates that the standard FLOPs regularization lacks semantic awareness, treating all tokens uniformly regardless of their informational value.
We further observe that tokens with higher activation scores tend to have higher IDF values, suggesting that IDF serves as an effective indicator of token importance within document representations. Motivated by this insight, we incorporate IDF as a guiding prior in the loss function to address the existing limitation.
IDF-aware Penalty.
Building on our observation, we introduce an IDF-aware Penalty to SPLADE-doc-distill by modifying the ranking objective. Specifically, we define an IDF-weighted ranking loss, $\mathcal{L}_{\text{rank-idf}}$, as a refinement of the original ranking loss $\mathcal{L}_{\text{rank}}$ in Eq. 1, in which the matching score in Eq. 2 is redefined as:
(5) $s(q,d_{i})=\sum_{t\in\mathcal{V}}\text{idf}(t)\cdot q_{t}\cdot d_{i,t},$
with $\text{idf}(t)$ denoting the IDF value of token $t$. This adjustment alters the gradient dynamics during training. The final training objective combines $\mathcal{L}_{\text{rank-idf}}$ with a FLOPS-based regularization term:
(6) $\mathcal{L}=\mathcal{L}_{\text{rank-idf}}+\lambda\cdot\mathcal{L}_{\text{FLOPS}}.$
For each token $t$, the gradient of the total loss with respect to its document-side activation $d_{i,t}$ can be decomposed as:
(7) $\frac{\partial\mathcal{L}}{\partial d_{i,t}}=\frac{\partial\mathcal{L}_{\text{rank-idf}}}{\partial d_{i,t}}+\lambda\cdot\frac{\partial\mathcal{L}_{\text{FLOPS}}}{\partial d_{i,t}}.$
The first term, originating from the ranking loss, is proportional to the token’s IDF value:
$\frac{\partial\mathcal{L}_{\text{rank-idf}}}{\partial d_{i,t}}\propto\text{idf}(t)\cdot q_{t}\cdot(\text{softmax}_{\text{stu}}-\text{softmax}_{\text{tea}}).$
while the FLOPS regularization imposes a uniform penalty (e.g., $2d_{i,t}$).
This composition reveals an important insight: For tokens with high $\text{idf}(t)$, the gradient from the ranking loss dominates, encouraging the model to preserve their activations. In contrast, tokens with low IDF values receive weaker ranking gradients and are thus more susceptible to sparsification by the FLOPS regularization. As a result, the model learns to retain informative tokens while effectively penalizing unimportant ones, leading to efficient retrieval with minimal performance degradation.
@@ 5.1.2. Model Training
For the teacher models of the pre-training phase, we employ SOTA dense and sparse retrievers , namely gte-large-en-v1.5555https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5 and opensearch-neural-sparse-encoding-v1666https://huggingface.co/opensearch-project/opensearch-neural-sparse-encoding-v1. The total number of pre-training steps is 150,000. For each step, we sample 48 queries and 8 hard samples(1 positive document and 7 hard negative documents) for every query. We set the learning rate to 5e-5 and the FLOPS $\lambda_{d}$ to 1e-7. The max input length is set to 128. The constant $S$ is set to 10 for scaling back the assembled scores. In the fine-tuning phase, we utilize an ensemble teacher model comprising the two siamese retrievers used in the pre-training phase, as well as two cross-encoder re-rankers [^7] We’re using https://huggingface.co/castorini/monot5-3b-msmarco-10k and https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-12-v2 . The total number of training steps is 50,000. For each step, we sample 40 queries and 10 hard negatives for each query. We set the learning rate to 2e-5 and the FLOPS $\lambda_{d}$ to 0.02. The max input length is set to 256. And the IDF values are calculated using the documents of MS MARCO dataset. The constant $S$ is set to 30 for scaling back the assembled scores.
@@ 5.4.1. Impact on search relevance
To assess the impact of IDF-aware penalty on the zero-shot search relevance, we conduct an ablation study employing different IDF settings on BEIR benchmark. In the default setting, we use fixed IDF values derived from MS MARCO dataset. We also conduct experiments using IDF values from the corresponding BEIR datasets. With different settings, we examined the impact of (1) employing IDF-aware penalty in training phase (2) retrieval with IDF values derived from different sources. We conduct these experiments on our model and SPLADE-doc-distill. The experiment results are shown in Table 5. From the experiment results, we obtain several conclusions: (1) The IDF-aware penalty boost the model search relevance at large margin. Training and inference with IDF, both our model and SPLADE-doc-distill achieve much better search relevance. (2) For models without pre-training phase, using IDF derived from the test set has better search relevance compared with fixed IDF derived from training data. However, if the model has undergone extensive pre-training on large-scale data using the fixed IDF, the conclusion is the opposite.
@@ 5.4.2. Impact on retrieval efficiency and index size
Experiments are conducted to measure the impact of IDF-aware penalty on retrieval efficiency. We conducted experiments to quantify the relationship between the expansion rate(average token number per expanded documents), retrieval efficiency(FLOPS number), and search relevance for our model. The comparison is made between our model with IDF-aware penalty and our model trained without IDF-aware penalty. The experimental results are depicted in Figure 4. The experimental findings demonstrate that the IDF-aware penalty substantially augments the retrieval efficiency. For models with similar expansion rate, the FLOPS number of those trained with IDF-aware penalty is much smaller.
Figure 4. The average FLOPS number, average token number per document and search relevance on BEIR datasets for models with and without IDF-aware penalty. The models are trained with different $\lambda_{d}$.
@@ 6. Conclusion
In this paper, we proposed two novel approaches to significantly improve the search relevance of inference-free learned sparse retrievers while maintaining high efficiency. We introduced IDF-aware penalty to mitigate the uniform penalty on tokens, boosting both relevance and efficiency. We also developed a heterogeneous ensemble knowledge distillation framework leveraging strong dense and sparse retrievers for pre-training, enhancing generalization. Extensive experiments validate our methods’ effectiveness. Our model achieves SOTA performance among inference-free sparse retrievers on BEIR, while maintaining end-to-end latency only 1.1x that of BM25.